OpenStudy (anonymous):

determine the limit of the sequence if convergent. An= (1+(2/n))^n

OpenStudy (anonymous):

First establish that \(a_n\) is a convergent sequence. Show that it's monotonic and bounded, then you can go ahead with finding the limit.

OpenStudy (anonymous):

SaG, I know thats the general form to go about, and im sure that ln is used, however i have zero to none on ideas to solve this problem. If you have specifics that would be sith tons of awesomeness

OpenStudy (anonymous):

To show it's monotonic: Consider the function \[f(x)=\left(1+\frac{2}{x}\right)^x\] You can check for monotonicity by examining the derivative: \[\begin{align*}\ln f(x)&=x\ln\left(1+\frac{2}{x}\right)\\\\\\ \frac{1}{f(x)}f'(x)&=\ln\left(1+\frac{2}{x}\right)+x\left(\frac{-\dfrac{2}{x^2}}{1+\dfrac{2}{x}}\right)\\\\\\ \frac{1}{f(x)}f'(x)&=\ln\left(1+\frac{2}{x}\right)-\frac{2}{x+2}\\ f'(x)&=\left(1+\frac{2}{x}\right)^x\left(\ln\left(1+\frac{2}{x}\right)-\frac{2}{x+2}\right) \end{align*}\] Find the critical points, etc. Figure out for which \(x\) the function is either strictly decreasing or increasing.

OpenStudy (anonymous):

The limit is \[ e^2\]

OpenStudy (anonymous):

Use L'Hospital Rule

OpenStudy (anonymous):

on log(f(x))

OpenStudy (anonymous):

To show it's bounded: Find an upper bound (if the sequence is increasing) or a lower bound (if the sequence is decreasing). Recalling the limit \[\lim_{x\to\infty}\left(1+\frac{1}{x}\right)^x=e\] might help. You could also try listing the first few terms to see if you can extrapolate what the limit might be.